If 1495 J of heat is needed to raise the temperature of a 325 g sample of a metal from 55.0°C to 66.0°C, what is the specific heat capacity of the metal?
1
Expert's answer
2019-03-11T08:20:06-0400

c = Q/mΔT

ΔT = 66-55=11 (K)

c = 1495 / (0.325*11) = 418.18 (J/kg*K)